# Programmable Cognitive Bias in Social Agents

### v2

@@ 7.3 Reproducibility
Bias Type Paradigm Category Metric Prompt RepE Lin. RepE Proj. Finetune
Authority Effect
Milgram
Obedience
Monotonicity NDCG 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
@@
Stanford
Prison
Monotonicity NDCG 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
@@
Bandwagon Effect
Asch’s
Line
Monotonicity NDCG 1.00 (0.00) 0.99 (0.01) 1.00 (0.01) 1.00 (0.00)
@@
Hotel
Towel
Monotonicity NDCG 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 0.99 (0.01)
@@
Confirmation Bias
Wason
Selection
Monotonicity NDCG 0.98 (0.02) 0.98 (0.04) 0.99 (0.01) 1.00 (0.00)
@@
Biased
Information
Monotonicity NDCG 0.97 (0.04) 1.00 (0.00) 0.98 (0.03) 1.00 (0.00)
@@
Framing Effect
Asian
Disease
Monotonicity NDCG 0.99 (0.00) 1.00 (0.00) 1.00 (0.00) 0.99 (0.01)
@@
Investment
/Insurance
Monotonicity NDCG 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 0.99 (0.01)
@@
Table 2: Comprehensive controllability results across four bias types, eight experimental paradigms, and four control methods. Values are presented as Mean (Standard Deviation), with the best highlighted in blue, and the worst in red (ties are resolved by mean, then standard deviation). Representation Engineering (RepE Linear and RepE Projection) consistently achieves the best balance of high monotonicity (NDCG $\approx 1.00$, Spearman’s $\rho\approx 0.9$) and superior smoothness (lowest $\Delta^{1},\Delta^{2}$ values), making it ideal for fine-grained, predictable control. Finetuning offers a robust alternative, while Prompt Numerical control provides the largest expressiveness range.
7.3.1 Across Models
Setup. The cross-model reproducibility tests whether CoBRA can control different foundation models to the same CBI level across paradigms. We calibrate Control Coefficients under one paradigm (e.g., Milgram Obedience paradigm (Milgram, 1963) for authority bias) on four models: Mistral, Llama, Qwen, and Deepseek. Then we evaluate their CBI difference in another paradigm (e.g., Stanford Prison paradigm (Haney et al., 1973) for authority bias) to test reproducibility. Lower difference indicates better reproducibility. We also include two reference baselines to better understand how well CoBRA can perform. The first baseline is No control, representing the natural CBI difference across all testing models in the target test paradigm. The second one is marked as Control, which represents the CBI difference when using identical natural language-based control (e.g., ”Pretend you have completely [bias type]”)
Results. In Fig. 8, the x-axis is the specified CBI in the calibrating paradigm, while the y-axis shows the CBI difference in the target paradigm. All CoBRA methods reduce cross-model reproducibility gaps below both baselines across most of the coefficient spectrum. RepE Projection attains the lowest mid-range gap (smooth internal trajectory).
Figure 8: Reproducibility of agent behavior across foundation models. The x-axis represents the Cognitive Bias Index in the calibrating paradigm, and the y-axis shows the measured Cognitive Bias Index difference (variances) across the models in the target test paradigm. All CoBRA methods significantly reduce this variance compared to baselines (No Control, Control). This means that CoBRA ensures different foundation models produce nearly identical behaviors for the same target bias level.
7.3.2 Across Sampling Temperatures
Setup. We test different control methods on the two authority bias experimental paradigms using the Mistral model across temperature $T\in\{0.1,0.2,...,1.0\}$. This experiment shows how temperature may affect CoBRA.
Results. From the Fig. 9, we can see the controlled result is insensitive to temperature changes. Representation engineering methods show an even smaller gap across temperatures. The results indicate that the Control Coefficient we obtained under one temperature can be directly applied in experiments using a different temperature, demonstrating the reproducibility across sampling temperatures.
Figure 9: Reproducibility experiments across sampling temperatures. The x-axis is the Control Coefficient. The y-axis is the resulting Cognitive Bias Index in the target bias paradigm. Each colored line represents a different temperature setting, ranging from a deterministic T=0.1 to a more random T=1.0. The fact that all the lines are tightly clustered and follow the same path demonstrates that the control is not affected by changes in temperature, ensuring that CoBRA ensures reproducibility results across different sampling temperatures.
7.3.3 Across Reasoning Modes
Setup. We choose Mistral as our testing foundation model and use the two authority bias experimental paradigms for experiment setup. We compare the results for two modes, Direct and Reasoning. In Direct mode, agents will give an immediate answer directly after reading the question; while in Reasoning mode, agents will first give a brief reasoning before answering the question. We use a temperature $T=0.7$ for both modes and set the reasoning token number to 128. Since the model will give various reasoning paths for the same question, we select the sampling number to be 8 (8 different reasoning paths for one question) to reduce the variance of the results.
Results. Cross-paradigm results (Fig. 10) show both methods preserve the linear correlation across different paradigms. For RepE methods, the Direct and Reasoning modes share almost identical understanding across different paradigms. Within a single paradigm, RepE Linear and Projection produce almost overlapping curves for two modes (Fig. 11) while the Prompt Numerical method doesn’t. The first result (Fig. 10) indicates that the Control Coefficient obtained using Direct mode calibration can be seamlessly used in paradigms requiring reasoning. The second result (Fig. 11) shows that RepE methods achieve superior consistent and continuous control of cognitive bias – even during the reasoning, the model coherently sticks to the codified cognitive bias without deviating from the pre-specified CBI level.
Figure 10: Reproducibility of different reasoning modes across paradigms. The x-axis represents the Cognitive Bias Index measured in the Milgram Obedience Paradigm, while the y-axis shows the Cognitive Bias Index using the same control coefficients in the Stanford Prison Paradigm with Direct mode and Reasoning mode. The Control Coefficients generalize across different paradigms for both modes.
Figure 11: The x-axis represents the Control Coefficient, and the y-axis shows the CBI under the Milgram Obedience Paradigm (Milgram, 1963). Within a single paradigm, Representation Engineering methods produce nearly identical behavior for both modes, indicating that the programmed bias is stable and consistent even during complex reasoning.
7.4 Controllability
We assess fine-grained controllability quality under a foundation model with full internal access and restricted API-only access. We try to find how good different control methods are in CoBRA in different paradigms by measuring smoothness, monotonicity, and expressiveness.
Setup. For open-source models, we evaluate all control methods across all paradigms and calculate the evaluation scores for each method. For API-based models, we exclusively use the Prompt Numerical method. Without access to internal logits for calculating CBI, we estimate probabilities using frequencies from 10 samples per question (at temperature $T=0.7$). This estimation remains unbiased, as we have demonstrated CoBRA’s invariance under temperature.
Results. In Fig. 12 and Table 2, among all methods, representation engineering-based approaches provide the most reliable control, with high monotonicity and smoothness (lower values indicate greater smoothness). The RepE Projection method achieves the smoothest control, while the RepE Linear method provides the most monotonic and consistent control. The finetuning method shows average performance across all metrics. The Prompt Numerical method, as expected, performs worst in monotonicity and smoothness, indicating instability in the input space for manipulating cognitive biases. While it offers the largest expressiveness, its discontinuous input space limits practical use, as it cannot specify arbitrary levels of control (e.g., distinguishing between 50% and 50.1% bias). However, it remains useful for broad control.
Figure 12: Control curves for four open-source foundation models (Deepseek, Llama, Mistral, Qwen) across four control methods and their four experimental paradigms. The x-axis represents the normalized Control Coefficient, and the y-axis is the resulting CBI. This figure illustrates straightforwardly how CoBRA’s methods control agent bias in a continuous and fine-grained way.
In Fig. 13 and Table 3, for models via API, the Prompt Numerical method offers practical control. Even with black-box API access, calibrated prompt numerical control yields reproducible, near-monotonic trajectories. Model differences suggest certain APIs (e.g., Claude) are more suited for continuous social simulation control.
Figure 13: CoBRA provides effective control even for models accessed via black-box APIs, where only the Prompt Numerical method is applicable. The x-axis represents the Control Coefficient, while the y-axis depicts the resulting CBI under the Milgram Obedience Paradigm (Milgram, 1963). The control curves are near-monotonic.
@@ 7.4 Controllability
gpt 0.996 0.980 0.132 0.178 2.32
mistral-n 0.998 0.870 0.131 0.134 2.44
claude 1.00 1.00 0.150 0.134 3.00
gemini 0.980 0.942 0.269 0.437 2.76
Table 3: Controllability results on models via APIs demonstrate CoBRA’s effectiveness.
@@
CoBRA delivers (i) consistent and generalizable Cognitive Bias Index (CBI), (ii) reproducible control across models, temperatures, and reasoning modes, and (iii) fine-grained control under both open-weight and API-only constraints. Method trade-offs: Prompt Numerical control offers the most expressive range, Representation Engineering (RepE) methods provide superior monotonicity and smoothness, making them ideal for fine-grained, predictable behavioral modulation. Finetuning serves as a robust, all-around alternative. Black-box results affirm external applicability without internal probability access. Collectively, these findings establish CoBRA as a reliable and versatile toolkit for specifying and controlling agent behavior in reproducible social science simulations.
8 Demonstration
In this section, we show how social scientists can use CoBRA to simulate the Emotional Contagion experiment (Kramer et al., 2014), where the participant’s social influence susceptibility (i.e., bandwagon effect bias) can be manipulated. The simulated experiment validates a classical social science theory: agents exposed to varying levels of negative posts exhibited predictable sentiment shifts based on their bias levels programmed by CoBRA, with higher Bandwagon Effect levels leading to stronger emotional contagion.
8.1 Methodology
@@
Experiment setup. To create a realistic social stimulus, we generate a corpus of 1,000 unique and emotionally negative Facebook-style posts using a suite of state-of-the-art LLMs (GPT-4.1, Claude-Sonnet-4, Gemini-2.5-Pro, and Grok-4). Each agent is exposed to a simulated feed where the number of negative posts (Independent Variable) is systematically varied from 0 to 15 by sampling from our corpus. After exposure, agents are prompted to generate their own post.
@@
Measurement: We quantify the emergent behavior of emotional contagion by computing a Sentiment Score for each generated post using the ‘twitter-roberta-base-sentiment-latest‘ model (Camacho-collados et al., 2022; Loureiro et al., 2022), defined as ($p_{pos}-p_{neg}$).
Figure 14: Baseline control with implicit natural language specification. The x-axis shows exposure to negative posts (the social stimulus), and the y-axis shows the resulting emotional contagion (the behavioral outcome). Agents are assigned bias settings from “No bias” to “Much bias,” but the resulting emotional contagion curves collapse onto one another: trajectories are largely overlapping, and their ordering is unstable. The absence of a clear separation structure shows that implicit natural language does not produce reliable, fine-grained control over the emergent social behavior.
Figure 15: Demonstration of CoBRA’s control. Agents are programmed with five distinct levels of the Bandwagon Effect (CBI values in legend). The x-axis shows exposure to negative posts (the social stimulus), and the y-axis shows the resulting emotional contagion (the behavioral outcome). The clear dose-response relationship confirms that controlling the underlying cognitive bias with CoBRA produces a predictable effect on a complex, emergent social behavior.
8.2 Results
@@
The results from the baseline method, shown in Figure 14, highlight the limitations of conventional approaches. The implicit natural language specification fails to produce reliable or fine-grained control. The emotional contagion curves for the different bias levels collapse onto one another, with largely overlapping trajectories and no stable ordering. This demonstrates that simple instructions are insufficient for precisely controlling complex cognitive dynamics.
In stark contrast, the results for CoBRA, presented in Figure 15, provide strong evidence for our validation goal. We observe a pronounced dose-response relationship: the degree of emotional contagion exhibited by the agents is strongly correlated with their pre-programmed level of the Bandwagon Effect. This predictable, fine-grained differentiation between agents confirms that the variable controlled by CoBRA has a direct and measurable impact on the emergent behavior.
This demonstration confirms that CoBRA’s application extends far beyond its Classic Social Experiment Testbeds. CoBRA is not merely manipulating surface-level responses. Instead, it successfully modulates an underlying cognitive process, which in turn governs how agents behave in novel and complex social simulations.
9 Limitation

### v3

@@ 7.2.1. Apparatus
Cognitive Bias Type Social Experimental Paradigm Category Metric Prompt Numerical RepE Linear RepE Projection Fine-tuning
Authority Effect Milgram Obedience (Milgram, 1963) Monotonicity NDCG 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
@@
Stanford Prison (Haney et al., 1973) Monotonicity NDCG 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
@@
Bandwagon Effect Asch’s Line (Asch, 1951) Monotonicity NDCG 1.00 (0.00) 0.99 (0.01) 1.00 (0.01) 1.00 (0.00)
@@
Hotel Towel (Goldstein et al., 2008) Monotonicity NDCG 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 0.99 (0.01)
@@
Confirmation Bias Wason Selection (Wason, 1960) Monotonicity NDCG 0.98 (0.02) 0.98 (0.04) 0.99 (0.01) 1.00 (0.00)
@@
Biased Information (Lord et al., 1979) Monotonicity NDCG 0.97 (0.04) 1.00 (0.00) 0.98 (0.03) 1.00 (0.00)
@@
Framing Effect Asian Disease (Tversky and Kahneman, 1981) Monotonicity NDCG 0.99 (0.00) 1.00 (0.00) 1.00 (0.00) 0.99 (0.01)
@@
Investment/Insurance (Kahneman and Tversky, 2013) Monotonicity NDCG 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 0.99 (0.01)
@@
Table 2. Comprehensive controllability results on open-source models across four bias types, eight experimental paradigms, and four control methods. Values are reported as Mean (Standard Deviation across models), with the best highlighted in blue and the worst in red (ties broken by mean, then standard deviation). Representation Engineering (RepE Linear and RepE Projection) consistently achieves the best balance of high monotonicity (NDCG $\approx 1.00$, Spearman’s $\rho\approx 0.9$) and superior smoothness (lowest $\Delta^{1}$ and $\Delta^{2}$ values), making it ideal for fine-grained control. Fine-tuning provides a robust alternative, while Prompt Numerical control offers the largest expressiveness range.
7.2.2. Monotonicity
We measured the Monotonicity using two metrics: Soft monotonicity, where Normalized Discounted Cumulative Gain (NDCG) (Järvelin and Kekäläinen, 2002) compares actual sequences to ideal monotonic ones, and Strict monotonicity, where (Spearman’s $\rho$) measures exact rank preservation (Spearman, 1961).
Results. Table 2 shows that, on open-source models, all three learning-based control methods achieve near-perfect soft monotonicity (overall NDCG $\approx 0.99$$1.00$) and high strict monotonicity (overall Spearman’s $\rho\approx 0.9$). Representation Engineering (RepE Linear and RepE Projection) generally yields the best rank preservation: averaged over all bias types and paradigms, RepE Linear attains the highest $\rho$ (0.94), with RepE Projection (0.89) and fine-tuning (0.91) close behind. Prompt Numerical control attains slightly lower $\rho$ (0.86), indicating that directly nudging the prompt is less reliable for preserving a strictly ordered progression of bias levels. The Control Baseline (Figure 10) behaves much less monotonically: $\rho<0.7$ for most scenarios. For closed-source API models, Table 3 indicates a similar picture in the Milgram Authority paradigm: all four models maintain high NDCG ($\geq 0.98$), and three (GPT-4o-mini, Claude, Gemini) reach Spearman’s $\rho$ comparable to or above the open-source average, with Claude almost perfectly monotone. This suggests that the same control scheme remains effective even when only sample-based probabilities are available.
7.2.3. Smoothness
We measured Smoothness using two related metrics. First-order smoothness captures local stability: it looks at how much the bias level changes, on average, when we move from one control setting to the next. If neighboring settings produce similar bias levels, first-order smoothness is high (i.e., changes are small). Second-order smoothness captures how consistent those changes are: it looks at how much the size of the step in bias level itself changes from one setting to the next. If the bias increases (or decreases) at a steady rate, second-order smoothness is high; if the rate keeps speeding up or slowing down, it is low.
Results. On open-source models (Table 2), Representation Engineering again provides the smoothest control. RepE Projection achieves the lowest overall first- and second-order differences ($\Delta^{1}=0.07$, $\Delta^{2}=0.04$), closely followed by RepE Linear ($\Delta^{1}=0.09$, $\Delta^{2}=0.04$). Fine-tuning is slightly less stable ($\Delta^{1}=0.10$, $\Delta^{2}=0.06$) but still produces relatively gradual changes in bias. In contrast, Prompt Numerical control yields noticeably larger discrete jumps (overall $\Delta^{1}=0.18$, $\Delta^{2}=0.22$), confirming that small changes in the prompt coefficient can produce irregular shifts in behavior. The Control Baseline is smoother ($\Delta^{1}=0.02$) due to its ineffective control. For closed-source models, Table 3 shows a qualitatively similar trend: smoothness varies across models, but at least one (Mistral-Nemo) achieves first- and second-order differences on par with, or better than, many open-source settings, indicating that our control procedure can still induce relatively stable transitions even when logits are unavailable and probabilities are estimated from samples.
7.2.4. Expressiveness
We measured Expressiveness by looking at how wide a range of bias levels the control can produce. In simple terms, it is the gap between the most biased behavior we observed and the least biased behavior we observed.
Results. For open-source models, Table 2 reveals a clear trade-off between expressiveness and regularity. Prompt Numerical control achieves the largest overall range (expressiveness =2.57=2.57), frequently spanning from strongly biased to strongly reverse-biased responses. RepE Linear, RepE Projection, and fine-tuning provide somewhat narrower but still substantial ranges (around 1.8–2.1), with RepE Projection often sacrificing a small amount of range in exchange for superior smoothness. The Control Baseline is not reliably expressive: mean expressiveness smaller than $0.26$. For closed-source API models, we observe similarly wide controllable ranges in the Milgram Authority paradigm (Table 3): Claude, in particular, attains very high expressiveness (3.000), with the other APIs also covering a range comparable to or exceeding that of most open-source settings. This suggests that our control framework generalizes well across both open-source and proprietary systems, even under black-box access.
@@ 7.2.4. Expressiveness
gpt 0.995 (0.003) 0.980 (0.060) 0.132 (0.018) 0.178 (0.042) 2.320 (0.086)
mistral-n 0.998 (0.001) 0.870 (0.064) 0.131 (0.006) 0.134 (0.015) 2.436 (0.067)
claude 1.000 (0.003) 1.000 (0.004) 0.150 (0.012) 0.134 (0.039) 3.000 (0.088)
gemini 0.980 (0.004) 0.942 (0.024) 0.269 (0.026) 0.437 (0.054) 2.760 (0.075)
Table 3. Controllability results on models via API across Milgram paradigms for Authority bias. Values show Mean (SE).
7.2.5. Generalization
We test whether Control Coefficients calibrated on the Investment/Insurance paradigm (Kahneman and Tversky, 2013) (individuals are more likely to purchase insurance when risks are described as potential losses rather than as missed gains) generalize to the Asian Disease paradigm (Tversky and Kahneman, 1981) (Participants are more risk-averse under positive frames and more risk-seeking under loss frames). We select 10 diverse agent personas from the persona pools in Generative Agents (Park et al., 2023), with varied specifications. For each persona, we sweep over all possible Control Coefficients. For each coefficient, we measure the resulting CBI in both the Investment/Insurance Paradigm and the Asian Disease Paradigm. This generates paired CBI sequences, $(y_{k}^{\text{invest}},y_{k}^{\text{asian}})$, for each persona across the two experiments. We also evaluate CoBRA using the Behavioral Regulation Engine, activating only one method at a time.
Refer to caption A line plot demonstrating cross-paradigm transferability of control coefficients for framing effects. The figure plots the Cognitive Bias Index measured in the Asian Disease paradigm against the index measured in the Investment and Insurance paradigm using the same control coefficients. Thin lines represent individual agent personas, a bold line shows the population average, and a shaded band indicates variability, illustrating that coefficient rankings are preserved across paradigms with high consistency.
Figure 11. Cross-paradigm transferability of Control Coefficients. We calibrate Control Coefficients on the Investment/Insurance Paradigm to produce varying levels of framing effect, then apply these same Coefficients to the Asian Disease Paradigm. The x-axis represents the Cognitive Bias Index (CBI) of framing effect measured in the Investment/Insurance Paradigm by varying Control Coefficients, while the y-axis shows the CBI measured in the Asian Disease Paradigm using the same Control Coefficients. Each thin line represents a unique agent persona; the bold line shows the population average; the shaded region indicates $\pm 1$ SE bands. Transferred Coefficients preserve ranking across paradigms (Spearman’s $\rho>0.96$, all $p<0.01$) with cross-persona consistency (CV $5\%$), demonstrating robust transferability.
Results. We visualizes the key results in Figure 11, which reveals two findings. First, rank preservation: Control Coefficients that produce stronger bias in the source paradigm also produce stronger bias in the target paradigm across all personas (Spearman’s $\rho>0.96$, all $p<0.01$). This monotonic relationship shows that the relative ordering of control strength transfers reliably. Second, cross-persona consistency: different personas exhibit nearly identical responses to the same transferred Control Coefficient (coefficient of variation $\text{CV}<5\%$), indicating that control effects generalize robustly across diverse agent characteristics. Together, these results suggest that CoBRAdoes not simply overfit to a particular text description: Control Coefficients learned on one paradigm transfer effectively to unseen paradigms without paradigm-specific recalibration. See Appendix D.2 for full statistical details.
@@ 7.2.5. Generalization
CoBRA provides (i) a generalizable Cognitive Bias Index (CBI), (ii) reproducible and fine-grained control across models, temperatures, and reasoning modes. In terms of method trade-offs, Prompt Numerical control offers the broadest expressiveness, while Representation Engineering (RepE) methods achieve superior monotonicity and smoothness, making them well-suited for precise, predictable behavioral modulation. Fine-tuning serves as a robust all-purpose alternative. Together, these findings establish CoBRA as a reliable toolkit for specifying and controlling agent behavior in social science simulations.
8. Demonstration
In this section, we show how social scientists can use CoBRA to simulate the Emotional Contagion experiment (Kramer et al., 2014), in which participants’ susceptibility to social influence (i.e., Bandwagon Effect bias) can be manipulated. Our simulated experiment aims to validate a classic social science finding: agents exposed to varying levels of negative posts exhibit predictable sentiment shifts based on the Bandwagon Effect levels programmed by CoBRA, with higher Bandwagon Effect leading to stronger emotional contagion.
(a) Baseline control with implicit natural language specification.
(b) CoBRA control. In the legend, ”coef.” denotes the control coefficient used to achieve the specified CBI.
Two plots comparing baseline control and CoBRA-based control of the bandwagon effect in an emotional contagion scenario. The up plot shows that implicit natural language specifications result in overlapping and unstable emotional contagion trajectories across bias settings. The down plot shows that CoBRA produces a clear, ordered dose–response relationship, where increasing control coefficients lead to predictable changes in emotional contagion as exposure to negative posts increases, with shaded regions indicating variability.
Figure 12. Comparison of baseline control and CoBRA control of the Bandwagon Effect in an emotional contagion scenario. The x-axis shows exposure to negative posts (social stimulus), and the y-axis shows emotional contagion (behavioral outcome). Shaded regions indicate $\pm 1$ SE bands.
@@ 8. Demonstration
Experiment setup. To create a realistic social stimulus, we generate a corpus of 1,000 unique, emotionally negative, Facebook-style posts using a suite of state-of-the-art LLMs (GPT-4.1, Claude-Sonnet-4, Gemini-2.5-Pro, and Grok-4). Each agent is exposed to a simulated feed where the number of negative posts (Independent Variable) is systematically varied from 0 to 15 by sampling from our corpus. After exposure, agents are prompted to generate their own post.
@@
Measurement. We quantify the emergent behavior of emotional contagion by computing a Sentiment Score for each generated post using the ‘twitter-roberta-base-sentiment-latest‘ model (Camacho-collados et al., 2022; Loureiro et al., 2022). A lower Sentiment Score means the posts are more negative, and a higher score means more positive.
@@
Results. The results from the baseline method (Fig. 12(a)) show that implicit natural language specifications fail to produce effective control: the emotional contagion curves for different intended bias levels largely collapse onto one another, with overlapping trajectories and no stable ordering. This indicates that simple instructions are insufficient for precisely controlling complex cognitive dynamics. In contrast, the CoBRA results (Fig. 12(b)) provide strong evidence for our validation goal. We observe a clear dose–response relationship: the degree of emotional contagion is strongly correlated with the agents’ pre-programmed bandwagon effect levels. This predictable, fine-grained separation across agents indicates that CoBRA controls a stable behavioral tendency rather than a specific prompt pattern.
This demonstration occurs in an open-ended task, rather than multiple-choice testbeds. This suggests that CoBRA’s effects are not limited to superficial changes on a single questionnaire, but instead reflect modulation of an underlying context‑sensitivity process that generalizes to more complex social simulations.
9. LIMITATIONS & FUTURE WORK